A body has a charge of2 micro coloumb . if it has 2.5multiplies 10power 3 protons then number of electrons?

Grade:12th Pass

1 Answers

Eshan
5 years ago
Dear student, Assuming the charge to be-2\mu C,

Total charge of protons-Total charge of electrons=-2\mu C
\implies n_pe-n_e e=-2\mu C
\implies n_e=n_p+\dfrac{2\times 10^{-6}}{1.6\times 10^{-19}}
\implies n_e\approx 1.25\times 10^{13}